# T HREAD : Thinking Deeper with Recursive Spawning

### v1

@@ Abstract
Large language models (LLMs) have shown impressive capabilities across diverse settings, but still struggle as the length and complexity of the context increases. To address this challenge, we propose Thinking Recursively and Dynamically (ThReaD). Thread frames model generation as a thread of execution that, based on the context, can run to completion or dynamically spawn new threads. By spawning, threads can offload work (e.g., thinking, retrieving information) to child threads, which only return tokens needed for the parent thread to do its work. In effect, this enables the model to adapt, as needed, the amount of intermediate work used to produce tokens. We apply Thread in the settings of LLM task solving and question answering, where the dynamic threading allows the model to recursively decompose the given task or question into progressively simpler sub-problems that can be solved by separate child threads. We test Thread, implemented using a few-shot learning approach, on diverse benchmarks for agent tasks and data-grounded question answering. Thread achieves state-of-the-art performance with GPT-4 and GPT-3.5 on these benchmarks, including ALFWorld, TextCraft, and WebShop, along with two new benchmarks, DataCommons QA and MIMIC-III ICU QA. In addition, Thread outperforms existing frameworks by 10% to 50% absolute points with smaller models, including Llama-3-8b and CodeLlama-7b.
@@ 1 Introduction
Large Language Models (LLMs) have shown success in diverse settings [34, 12], but their performance degrades as context length and complexity grows [4, 19, 26]. This constraint limits their efficacy in settings that require more work (e.g., thinking, retrieving information, analyzing, etc.) than can fit into a concise line of generation. To address this limitation, we propose Thinking Recursively and Dynamically (ThReaD).
@@
Thread improves the flexibility of model generation, allowing the model to adapt, through recursive spawning, the amount of work it does based on the given problem, without overextending its context. Importantly, this unfolds with limited need for explicit rules or hard coded logic. Instead, Thread relies on the model’s ability to infer the appropriate continuation given the context of each thread at the time of execution. Finally, Thread is agnostic to the type of sequence. Depending on the underlying model, Thread can be applied in varied settings, including multi-modal applications, and fulfill varied purposes (e.g., writing programs, carrying out calculations, augmenting data, generating thoughts, retrieving information, interacting with an environment, robot manipulation, etc.).
@@
Figure 1: Thread with join synchronization. Thread frames model generation as an execution thread that can dynamically spawn new threads. In the above example with join synchronization, when a thread spawns a child, it pauses generation until feedback is returned. Child threads generate starting from context that derives from their parent’s token sequence. When the child completes, it returns output tokens (colored bars), which are added to the context of the parent before it continues generating. $\phi$ and $\psi$ are functions that control information flow from parent to child and from child to parent, respectively, by defining the tokens that are propagated based on the thread’s token sequence.
@@
In this paper, we consider Thread in the settings of LLM task completion (Figure 2a) and question answering (Figure 2b). In these settings, Thread enables the LLM to recursively decompose a task, or question, into progressively simpler sub-problems that can be solved by child threads in a compartmentalized manner. A child thread infers the specifications of its sub-problem from its parent’s token sequence. If needed, child threads can troubleshoot their sub-problem, and spawn their own threads, all without distracting their parent. We test Thread, using a few-shot learning approach, on benchmarks consisting of question-answering and agent tasks. For these problems, we apply Thread with join synchronization, allowing parent threads to see feedback from child threads before defining their next step. For example, in the agent-based setting, if the model is given the task of cleaning a bowl and putting it on the countertop, the main thread may spawn a child with the sub-task of finding a bowl. This child can then spawn further threads for handling the complexities of navigating to and checking different locations for the bowl, with each thread returning only the information needed for its parent to proceed with its sub-task. Based on the feedback with regard to finding the bowl, the main thread can spawn a new thread to re-attempt the sub-task or spawn threads to execute the remaining sub-tasks (i.e., washing the bowl, putting the bowl on the countertop).
@@
Thread has several advantages over existing frameworks in these settings. To start, Thread addresses limitations of methods that require the model to solve the problem in one line of context [37, 38, 27, 29] by allowing the model to dynamically offload work by spawning child threads as described above. Further, unlike methods for task decomposition [17, 30, 25, 33, 32], Thread enables the model to adapt its decision-making in real time as it receives feedback from each step, as depicted in Figure 2. Existing methods have the model generate a plan at the start of the response and either do not allow for plan adaptation [17, 30, 32] or only adapt plans by adding more sub-steps within the existing plan upon task failure [25, 33]. Finally, Thread provides a more unified framework compared to these methods, which require separate prompting mechanisms for separate planner and executor modules [25, 33]. Instead, Thread can be implemented with the same few-shot prompt used for every thread at every step of the task completion or question answering.
@@
We evaluate Thread on benchmarks for agent tasks and data-grounded question answering [28, 36, 25, 10, 14]. Thread significantly outperforms prior methods, achieving state-of-the-art performance with GPT-4 and GPT-3.5 on ALFWorld, TextCraft, and WebShop, along with two new benchmarks, DataCommons QA and MIMIC-III ICU QA. In addition, Thread shows success with smaller models, outperforming prior methods by 10% to 50% absolute points across the benchmarks with Llama-3-8b and CodeLlama-7b.
@@ 2 Thread framework
Thread frames model generation as an execution thread that, given the context, can run to completion or spawn new threads. A thread consists of generating with a model, $G$, given a specific starting context, $c$. For the main thread, $c$ is the initial seed context, $c_{\_}{0}$ (e.g., context provided by a user). For each child thread, $c$ derives from the parent’s token sequence. A thread continues until it meets some termination criteria, such as generating a specific stop token.
2.1 Thread with join synchronization
We present Thread in a setting where, analogous to how the join() method is used in multithreaded programming, a parent thread waits for the execution of the child to complete before proceeding. When the child execution completes, the child’s output is returned directly to the parent. In this setting, Thread can be implemented using the recursive function shown in Algorithm 1. The function takes two inputs: the context for the thread, $c$, and the tokens generated so far by the thread, $Y$. We treat a token sequence (e.g., $c$ and $Y$) as a list of tokens.
@@ 2.1 Thread with join synchronization
Thread begins with $c=c_{\_}{0}$ and $Y=[\,]$. The model generates, $G(c+Y)$, one token at a time conditioning on the given context, $c$, appended with the growing sequence of tokens generated by the thread, $Y$.
If $Y$ spawns a child thread, then a new thread is created with Thread($\phi(Y),[\,]$ ) where $Y$ for the child thread is initialized as an empty list and $c$ is based on the token sequence of the parent, $\phi(Y)$. The output tokens of the child thread are appended to the parent’s token sequence and the parent continues generating. If $Y$ ends the thread, then the thread returns $Y$ (its full token sequence).
@@
2.2 Thread with alternative synchronization and spawning mechanisms
@@ 2.3 Defining parent-child information exchange with $\phi$ and $\psi$
The functions $\phi$ and $\psi$ control the propagation of information from parent to child and from child to parent, respectively. Similar to the synchronization and spawning mechanisms mentioned above, these functions can vary based on the setting in which Thread is implemented. For example, $\phi$ and $\psi$ can range from simple functions that compress or decompress information in the token sequence to entirely separate models that transform the sequence in more complex ways. In the section below, we describe how we define $\phi$ and $\psi$ when implementing Thread in the settings of LLM task completion and question answering.
@@ 2.4 Dynamically adapting intermediate work in token generation
The Thread framework enables a thread of model generation to spawn a child thread to produce output tokens that can serve as future tokens for the parent. In effect, this enables the model to adapt, as needed, the amount of work or computational steps used to produce different tokens. The intermediate work added through recursive spawning can range from deeper thinking and reasoning to interacting with external environments or sources of information. As depicted in Figure 3, the work associated with each part of the parent’s token sequence is represented by a thread tree organized based on connections between parent and child threads. These trees reflect how threads interact to form output tokens and the computational work associated with each component of the overall model generation.
3 Applying Thread for LLM task completion and question answering
In this paper, we test Thread in the settings of question answering and task completion with LLMs, where Thread enables the model to dynamically decompose the given problem into simpler sub-problems that are completed by separate threads. We apply Thread using an in-context learning approach, where the same few-shot prompt is used for every thread at every step of the task completion or question answering. Since these problems benefit from the model adapting to feedback as it defines its next steps, we apply Thread with the join synchronization described above.
Thread spawning and termination based on special stop tokens. We implement a spawning mechanism using a special stop token, $\omega_{\_}{listen}$, which pauses the thread generation and marks the start of the output the thread expects from the child. The context for the child thread is defined by $\phi$ based on the token sequence of the parent that occurs before $\omega_{\_}{listen}$. A thread ends when it generates the end token, $\omega_{\_}{end}$. As exemplified in Figure 2, we use => as $\omega_{\_}{listen}$ and END as $\omega_{\_}{end}$.
Implementing Thread with few-shot learning. We leverage a few-shot learning approach to implement Thread. The few-shot examples are comprised in a single prompt with examples of successful spawning and problem solving at different thread depths, including how the token $\omega_{\_}{listen}$ should be used. As a result, the approach does not require explicit rules to define when new threads should be spawned and how parent threads should respond to feedback from a child. Instead, these mechanisms are all implied by the examples provided in the few-shot prompt. Changing the spawning and parent-child dynamics simply requires modifying the prompt. All threads are provided the same few-shot prompt for every step of the problem completion. Thus, Thread only requires generating a single prompt for a given setting. This prompt, $q$, is prepended to the context, $c=q+c$, for each thread, forming the full context from which the thread will generate.
@@ 3 Applying Thread for LLM task completion and question answering
Threads performing actions in an environment. Threads can listen for feedback when performing an action in an environment the same way they listen for feedback from a child thread. They simply generate tokens that represent the action and then produce the token $\omega_{\_}{listen}$ to listen for feedback from the environment. The action, $a$, is then executed in the environment, $E$, and the output, $o=E(a)$, is appended to the thread’s token sequence the same way it is done for the feedback from a child thread (Algorithm 2). The thread can then continue generating based on the feedback.
Error handling. We do not implement any explicit mechanisms for error handling and instead rely on the inherent reasoning ability of LLMs to respond to unexpected feedback from child threads or from the environment.
@@ 4 Experiments
We evaluate Thread on 5 benchmarks for LLM agent task completion and question answering using large and small models, including GPT-4, GPT-3.5, Llama-3-8b, Llama-2-7b, and CodeLlama-7b. For each benchmark, we use the same prompt for all models.
Thread significantly outperforms prior methods, achieving state-of-the-art performance with GPT-3.5 and GPT-4 on ALFWorld, TextCraft, WebShop, DataCommons QA, and MIMIC-III ICU QA. In addition, Thread shows success with smaller models, outperforming prior methods by 10% to 50% absolute points across the benchmarks with Llama-3-8b and CodeLlama-7b.
We include details regarding models, prompts, and experiments (including ablations) in the appendix. We release all code and data at https://github.com/philipmit/thread.
@@ 4.1 ALFWorld
ALFWorld is a suite of text-based environments, implemented in TextWorld, designed to align with the embodied ALFRED benchmark [28]. It includes 6 different task types that instruct the agent to accomplish a goal (e.g., wash a bowl and put it on countertop) by interacting with a simulated household with actions defined in text. Following Yao et al., [37], we evaluate an agent on 134 unseen evaluation games, including six task types: Pick, Clean, Heat, Cool, Look, and Pick2. Like previous methods, such as ReAct [37], we implement Thread with one few-shot prompt per task. We also test Thread using task-general prompting.
Table 1: ALFWorld task-specific success rates (%). Results are separated based on whether the method requires the agent to have access to external memory constructed from previous experiences with the task. *As reported in Kagaya et al., [15] (RAP) and Fu et al., [5] (AutoGuide).
Model Requires ext. mem. Method All Pick Clean Heat Cool Look Pick2
GPT-3.5 Yes Reflexion [27] 76.1 75.0 80.6 69.6 76.2 83.3 70.6
AdaPlanner [29] 82.8 91.7 87.1 82.6 95.2 50.0 82.4
RAP* [15] 85.8 95.8 87.1 78.3 90.5 88.9 70.6
AutoGuide* [5] 79.1 - - - - - -
No ReAct [37] 53.7 45.8 48.4 69.6 66.7 55.6 35.3
ADaPT [25] 82.1 87.5 83.9 78.3 90.5 83.3 64.7
@@
GPT-4 Yes RAP* [15] 94.8 95.8 90.3 100 95.2 100 88.2
No ReAct [37] 87.3 83.3 77.4 95.7 85.7 100 88.2
ADaPT [25] 91.0 91.7 87.1 95.7 90.5 94.4 88.2
@@
Requires ext. mem. Method Llama-3-8b Llama-2-7b CodeLlama-7b
Yes Reflexion [27] 25.4 11.2 27.6
AdaPlanner [29] 28.4 11.9 29.9
No ReAct [37] 20.1 12.7 23.1
ADaPT [25] 30.6 15.7 35.1
@@
Model All Pick Clean Heat Cool Look Pick2
GPT-3.5 95.5 100 96.8 82.6 95.2 100 100
Llama-3-8b 49.3 58.3 38.7 56.5 57.1 72.2 11.8
CodeLlama-7b 61.9 41.7 87.1 65.2 61.9 38.9 64.7

### v2

@@ Abstract
Large language models (LLMs) have shown impressive capabilities across diverse settings, but still struggle as the length and complexity of the context increases. To address this challenge, we propose Thinking Recursively and Dynamically (ThReaD). Thread frames model generation as a thread of execution that, based on the context, can run to completion or dynamically spawn new threads. By spawning, threads can offload work (e.g., thinking, retrieving information) to child threads, which only return tokens needed for the parent thread to do its work. We apply Thread in the settings of LLM task solving and question answering, where the dynamic threading allows the model to recursively decompose the given task or question into progressively simpler sub-problems that can be solved by separate child threads. We test Thread, implemented using a few-shot learning approach, on diverse benchmarks for agent tasks and data-grounded question answering. Thread achieves state-of-the-art performance with GPT-4 and GPT-3.5 on these benchmarks, including ALFWorld, TextCraft, and WebShop, along with two new benchmarks, DataCommons QA and MIMIC-III ICU QA. In addition, Thread outperforms existing frameworks by 10% to 50% absolute points with smaller models, including Llama-3-8b and CodeLlama-7b.
T HREAD : Thinking Deeper with Recursive Spawning
@@ 1 Introduction
footnotetext: Correspondence to Philip Schroeder at pschro@mit.edu. Source code is available at https://github.com/philipmit/thread.
Large Language Models (LLMs) have shown success in diverse settings Wei et al. (2022); Huang and Chang (2023), but their performance degrades as context length and complexity grows Dziri et al. (2023); Liu et al. (2024); Qin et al. (2023). This constraint limits their efficacy in settings that require more work (thinking, retrieving information, analyzing, etc.) than can fit into a concise line of generation. To address this limitation, we propose Thinking Recursively and Dynamically (ThReaD).
@@
Figure 1: Thread with join synchronization. Thread frames model generation as an execution thread that can dynamically spawn new threads. In the example with join synchronization, when a thread spawns a child, it pauses generation until feedback is returned. Child threads generate starting from context derived from their parent’s token sequence. When the child completes, it returns output tokens (colored bars), which are added to the context of the parent before it continues generating. $\phi$ and $\psi$ are functions that control information flow from parent to child and from child to parent, respectively, by defining the tokens that are propagated based on the thread’s token sequence.
Thread improves the flexibility of model generation, allowing the model to adapt, through recursive spawning, the amount of work it does based on the given problem, without overextending its context. Importantly, this unfolds with limited need for explicit rules or hard coded logic. Instead, Thread relies on the model’s ability to infer the appropriate continuation given the context of each thread at the time of execution. Finally, Thread is agnostic to the type of sequence. Depending on the underlying model, Thread can be applied in varied settings, including multi-modal applications, and fulfill varied purposes (e.g., carrying out calculations, generating thoughts, retrieving information, robot manipulation).
@@
In this paper, we consider Thread in the settings of agent tasks (Figure 2a) and question answering (Figure 2b). In these settings, Thread enables the LLM to recursively decompose a task, or question, into progressively simpler sub-problems that can be solved by child threads in a compartmentalized manner. A child thread infers the specifications of its sub-problem from its parent’s token sequence. If needed, child threads can troubleshoot their sub-problem, and spawn their own threads, all without distracting their parent. We test Thread, using a few-shot learning approach, on benchmarks consisting of question-answering and agent tasks. For these problems, we apply Thread with join synchronization, allowing parent threads to see feedback from child threads before defining their next step. For example, in the agent setting, if the model is given the task of cleaning a bowl and putting it on the countertop, the main thread may spawn a child with the sub-task of finding a bowl. This child can then spawn further threads for handling the complexities of navigating to and checking different locations for the bowl, with each thread returning only the information needed for its parent to proceed with its sub-task. Based on the feedback with regard to finding the bowl, the main thread can spawn a new thread to re-attempt the sub-task or spawn threads to execute the remaining sub-tasks (i.e., washing the bowl, putting the bowl on the countertop).
@@
Thread has several advantages over existing frameworks in these settings. To start, Thread addresses limitations of methods that require the model to solve the problem in one line of context Yao et al. (2023); Shinn et al. (2024); Sun et al. (2023a) by allowing the model to dynamically offload work by spawning child threads. Further, unlike methods for task decomposition Khot et al. (2023); Sun et al. (2023b); Prasad et al. (2023); Wang et al. (2024, 2023), Thread enables the model to adapt its decision-making in real time as it receives feedback from each step, as depicted in Figure 2. Existing methods either do not allow for plan adaptation Sun et al. (2023b); Wang et al. (2023) or only adapt plans by calling pre-existing sub-task handlers Khot et al. (2023) or by adding more sub-steps within the existing plan when a sub-task fails Prasad et al. (2023); Wang et al. (2024). Finally, Thread provides a more unified framework compared to these methods, which require separate prompting mechanisms for separate planner and executor modules. Instead, Thread can be implemented with the same few-shot prompt used for every thread at every step of the task completion.
@@
We evaluate Thread on agent tasks and data-grounded question answering. Thread significantly outperforms prior methods, achieving state-of-the-art performance with GPT-4 and GPT-3.5 on ALFWorld, TextCraft, and WebShop, along with two new benchmarks, DataCommons QA and MIMIC-III ICU QA. Thread also shows success with smaller models, outperforming prior methods by 10% to 50% absolute points across the benchmarks with Llama-3-8b and CodeLlama-7b.
@@ 2 Thread Framework
Thread frames model generation as an execution thread that, given the context, can run to completion or spawn new threads. A thread consists of generating with a model, $G$, given a specific starting context, $c$. For the main thread, $c$ is the initial seed context, $c_{0}$ (e.g., context provided by a user). For each child thread, $c$ derives from the parent’s token sequence. A thread continues until it meets some termination criteria, such as a stop token.
@@
2.1 Thread with Join Synchronization
We present Thread in a setting where, analogous to how the join() method is used in multithreaded programming, a parent thread waits for the execution of the child to complete before proceeding. When the child execution completes, the child’s output is returned directly to the parent. In this setting, Thread can be implemented using the recursive function shown in Algorithm 1. The function takes two inputs: the context for the thread, $c$, and the tokens generated so far by the thread, $Y$. We treat a token sequence as a list of tokens.
Thread begins with $c=c_{0}$ and $Y=[\,]$. The model generates, $G(c+Y)$, one token at a time conditioning on the given context, $c$, appended with the growing sequence of generated tokens, $Y$.
[…]
@@ 2.1 Thread with Join Synchronization
2.2 Alternative Synchronization and Spawning Mechanisms
@@ 2.3 Defining Parent-Child Information Exchange with $\phi$ and $\psi$
The functions $\phi$ and $\psi$ control the propagation of information from parent to child and from child to parent, respectively. Similar to the synchronization and spawning mechanisms mentioned above, these functions can vary based on the setting in which Thread is implemented. For example, $\phi$ and $\psi$ can range from simple functions that compress or decompress information in the token sequence to entirely separate models that transform the sequence in more complex ways. In the section below, we describe how we define $\phi$ and $\psi$ when implementing Thread in the settings of agent tasks and question answering.
@@ 2.4 Dynamically Adapting Intermediate Work in Token Generation
The Thread framework enables a thread of model generation to spawn a child thread to produce output tokens that can serve as future tokens for the parent. In effect, this enables the model to adapt, as needed, the amount of work or computational steps used to produce different tokens. The intermediate work added through recursive spawning can range from deeper thinking and reasoning to interacting with external environments or sources of information. As depicted in Figure 3, the work associated with each part of the parent’s token sequence is represented by a thread tree organized based on connections between threads. These trees reflect how threads interact to form output tokens and the computational work associated with each component of the overall model generation.
3 Applying Thread for Agent Tasks and Question Answering
In this paper, we test Thread in the settings of question answering and agent tasks with LLMs, where Thread enables the model to dynamically decompose the given problem into simpler sub-problems that are completed by separate threads. We apply Thread using an in-context learning approach, where the same few-shot prompt is used for every thread at every step of the task completion or question answering. Since these problems benefit from the model adapting to feedback as it defines its next steps, we apply Thread with the join synchronization described above.
Thread spawning and termination based on special stop tokens. We implement a spawning mechanism using a special stop token, $\omega_{listen}$, which pauses the thread generation and marks the start of the output the thread expects from the child. The context for the child thread is defined by $\phi$ based on the token sequence of the parent that occurs before $\omega_{listen}$. A thread ends when it generates the end token, $\omega_{end}$. As exemplified in Figure 2, we use => as $\omega_{listen}$ and END as $\omega_{end}$.
Implementing Thread with few-shot learning. We leverage a few-shot learning approach to implement Thread. The few-shot examples are comprised in a single prompt with examples of successful spawning and problem solving at different thread depths, including how the token $\omega_{listen}$ should be used. As a result, the approach does not require explicit rules to define when new threads should be spawned and how parent threads should respond to feedback from a child. Instead, these mechanisms are all implied by the examples provided in the few-shot prompt. Changing the spawning and parent-child dynamics simply requires modifying the prompt. All threads are provided the same few-shot prompt for every step of the problem completion. Thus, Thread only requires generating a single prompt for a given setting. This prompt, $q$, is prepended to the context for each thread, $G(q+c+Y)$, forming the full context from which the thread will generate.
@@ 3 Applying Thread for Agent Tasks and Question Answering
Threads performing actions in an environment. Threads can listen for feedback when performing an action in an environment the same way they listen for feedback from a child thread. They simply generate tokens that represent the action and then produce the token $\omega_{listen}$ to listen for feedback from the environment. The action, $a$, is then executed in the environment, $E$, and the output, $o=E(a)$, is appended to the thread’s token sequence the same way it is done for the feedback from a child thread (Algorithm 2). The thread can then continue generating based on the feedback.
@@ 4 Experiments
We evaluate Thread on 5 benchmarks for agent tasks and question answering using large and small models, including GPT-4, GPT-3.5, Llama-3-8b, Llama-2-7b, and CodeLlama-7b. For each benchmark, we use the same prompt for all models.
We include details regarding models, prompts, and experiments (including ablations) in the appendix. Examples of the prompts are shown in Appendix  G. Details regarding the QA benchmarks are provided in Appendix  F. We release all code and data at https://github.com/philipmit/thread.
@@ 4.1 ALFWorld
ALFWorld is a suite of text-based environments, implemented in TextWorld, designed to align with the embodied ALFRED benchmark Shridhar et al. (2021). It includes 6 different task types that instruct the agent to accomplish a goal by interacting with a simulated household with actions defined in text. Following Yao et al. (2023), we evaluate an agent on 134 unseen evaluation games, including six task types: Pick, Clean, Heat, Cool, Look, and Pick2. Like previous methods, such as ReAct Yao et al. (2023), we implement Thread with one few-shot prompt per task. We also test Thread using task-general prompting.
Results.
Thread significantly outperforms all prior methods (Tables 1 and 2), including those that require the agent to have access to external memory consisting of past experiences with the task. With GPT-4, Thread achieves a combined success rate of 98.5% with 100% success in 4 of the 6 tasks. With GPT-3.5, Thread shows a combined success rate of 95.5%, outperforming all methods by over 9% absolute points. With smaller models (Llama-3-8b and CodeLlama-7b), Thread improves upon prior methods by 30% to 55% points.
Finally, Table 8 shows the results when testing Thread with task-general prompting. To implement the task-general prompt, we split the prompt into one set of examples for the main thread and a second set of examples for all other threads (with the same sets of examples used for all tasks). We see that GPT-3.5 achieves the same combined success rate with the task-general prompting as it does with the task-specific prompting. Further, while the performance of Llama-3-8b and CodeLlama-7b degrades with task-general prompting, it remains a significant improvement over task-specific prompting with prior methods (Table 2).
4.2 TextCraft
TextCraft is a text-based environment inspired by the crafting component of Minecraft Prasad et al. (2023). The tasks involve building Minecraft items with crafting commands using available resources from the environment. Following the work of Prasad et al. (2023), we evaluate Thread on the test set containing 200 samples. We implement Thread using a single few-shot prompt and compare its performance to previous methods used for TextCraft Prasad et al. (2023); Wang et al. (2024); Shinn et al. (2024); Yao et al. (2023).
Table 1: ALFWorld task-specific success rates (%). Results are separated based on whether the method requires the agent to access to external memory. *As reported in Kagaya et al. (2024) (RAP) and Fu et al. (2024) (AutoGuide).
Model
Requires
ext. mem.
Method All Pick Clean Heat Cool Look Pick2
GPT-3.5 Yes Reflexion Shinn et al. (2024) 76.1 75.0 80.6 69.6 76.2 83.3 70.6
AdaPlanner Sun et al. (2023a) 82.8 91.7 87.1 82.6 95.2 50.0 82.4
RAP* Kagaya et al. (2024) 85.8 95.8 87.1 78.3 90.5 88.9 70.6
AutoGuide* Fu et al. (2024) 79.1 - - - - - -
No ReAct Yao et al. (2023) 53.7 45.8 48.4 69.6 66.7 55.6 35.3
DecomP Khot et al. (2023) 84.3 91.7 87.1 82.6 90.5 83.3 64.7
ADaPT Prasad et al. (2023) 82.1 87.5 83.9 78.3 90.5 83.3 64.7
@@ 4.2 TextCraft
GPT-4 Yes RAP* Kagaya et al. (2024) 94.8 95.8 90.3 100 95.2 100 88.2
No ReAct Yao et al. (2023) 87.3 83.3 77.4 95.7 85.7 100 88.2
DecomP Khot et al. (2023) 89.6 87.5 87.1 91.3 90.5 94.4 88.2
ADaPT Prasad et al. (2023) 91.0 91.7 87.1 95.7 90.5 94.4 88.2
@@
Requires
ext. mem.
Method Llama-3-8b Llama-2-7b CodeLlama-7b
Yes Reflexion Shinn et al. (2024) 25.4 11.2 27.6
AdaPlanner Sun et al. (2023a) 28.4 11.9 29.9
No ReAct Yao et al. (2023) 20.1 12.7 23.1
DecomP Khot et al. (2023) 37.3 14.2 33.6
ADaPT Prasad et al. (2023) 30.6 15.7 35.1